OpenStudy (anonymous):

SIMPLIFY 1/5 times square root of 5n^3 times square root of 20n^5

OpenStudy (jmark):

= 1/5 (√5n^3) (√5*4 n^5) = 1/5 .√(5n^3) .√(5 2 n^5) = 1/5. √(5)^2 √n^3 . 2 .√n^5 = 2.√ n^2 . √ n^5 ICSE class 7 maths question paper - http://icse.edurite.com/icse-question-papers/icse-class-7-maths-question-papers.html